Astrid1:

@aZ @sNoWfLaKe0531 ss below

Astrid1:

snowflake0531:

Square them bro 85 = p^2 + 68

snowflake0531:

So p^2 = 17 So p = ...?

Astrid1:

1?

snowflake0531:

astrido._. p^2 = 17 would what p equal e.e

Astrid1:

17 .).

snowflake0531:

e.e so a^2 + b^2 =c^2 \[\sqrt68^2 + p^2 = \sqrt85^2\] which we can simplify to \[p^2+68=85\] then\[p^2 = 17\] then\[\sqrt p^2 =\sqrt17\] so \[p=\sqrt17\]